A long, straight wire carrying a current of 305 A is placed in a uniform magnetic field that has a magnitude of 6.78 × 10-3 T. The wire is perpendicular to the field. Find a point in space where the net magnetic field is zero. Locate this point by specifying its perpendicular distance from the wire.
Answer #1

The field will =0 where the B produced by the wire = the B field

Use B= (u0*I)/(2pi*R)
B = 6.78E-3T
6.78E-3=(u0*305)/(2pi*R)


R=.008997 m